I need help solving this problem 56-2(9+3*2)-9=. Please show me the steps to get to the answer.

Thanks!
Teisha

Answer by edjones(8007)

56-2(9+3*2)-9
=56-2(9+6)-9
=56-2*15-9
=56-30-9
=17
.
Ed